OpenStudy (anonymous):

find the fifth term in the sentence tn=12-2n

OpenStudy (anonymous):

anyone?

OpenStudy (jordanloveangel):

t5 = 12 - 2(5) = 12 - 10 = 2

OpenStudy (anonymous):

what are terms though

OpenStudy (jordanloveangel):

one 1 sec

OpenStudy (jordanloveangel):

second term tn+1 = 12 - 2(n+1) = 10 - 2n third term 10 - 2(n+1) = 8 - 2n fourth term 6 - 2n fifith term 4 - 2n

OpenStudy (jordanloveangel):

. once again, plug 2 into N. 6-2=4. 4^2=16 (12 - 2n) * 5 = 60 - 10n

OpenStudy (solomonzelman):

Nooo.

OpenStudy (jordanloveangel):

oh i did something wrong?

OpenStudy (solomonzelman):

We don't need 2nd 3rd and 4th term to find the fifth... \(\Large\color{blue}{ t_n=12-2n}\) is the formula for any term in a sequence, as I can see, so plug in t=5.

OpenStudy (solomonzelman):

you did it right at first, but I am not sure why you found the 1st and 2nd terms after that...

OpenStudy (solomonzelman):

the slope, or the difference, is -2.